# Approximating the coefficients of the Bessel functions

### v4

@@ 1.4. Weak convergence to the free convolution
Suppose $\geq 1, N\mathcal{R}\subset\mathbb{R}^{N}$ is a finite root system, and $\theta\in\theta(\mathcal{R})$ is nonnegative. Suppose $a\in\mathbb{R}^{N}$. There exists a unique Borel probability measure $\mu_{a}$ whose support is contained in the convex hull of $H(\mathcal{R})a$ such that
$E_{a}^{\mathcal{R}}(x)=\int_{\mathbb{R}^{N}}e^{\sum_{i=1}^{N}x_{i}\epsilon_{i}}d\mu_{a}(\epsilon)$
for all $x\in\mathbb{C}^{N}$. Furthermore, $\text{supp}(\mu_{a})\cap H(\mathcal{R})a$ is nonempty and the Borel probability measure
$\mu_{a}^{\text{sym}}\triangleq\frac{1}{|H(\mathcal{R})|}\sum_{h\in H(\mathcal{R})}\mu_{ha}$
@@
$J^{\mathcal{R}(\theta)}_{a}(x)=\int_{\mathbb{R}^{N}}e^{\sum_{i=1}^{N}x_{i}\epsilon_{i}}d\mu_{a}^{\text{sym}}(\epsilon)$
@@
From [dunkloperators, dunklbessel], if $\theta\in\theta(\mathcal{R})$ such that $\theta(r)$ has nonnegative real part for all $r\in\mathcal{R}$, then $\theta\in\Theta(\mathcal{R})$, or equivalently, $\mathcal{D}(\mathcal{R}(\theta))$ is invertible. See Subsection˜2.1 for the definitions of these notions. The invertibility of the Dunkl operator implies the existence of it associated eigenfunctions, see Theorems˜2.5 and 2.8. In particular, this implies that the eigenfunctions exist in the context of Theorem˜1.4.
We state a well-known conjecture related to an integral representation of the product of two Bessel functions; it can easily be extended to the product of an arbitrary number of Bessel functions. The conjecture also assumes that the multiplicity function is nonnegative.
@@
The measure $\mu_{a_{1},a_{2}}^{\mathcal{R}(\theta)}$ is clearly not unique, by the symmetry of $J_{a}^{\mathcal{R}(\theta)}(x)$. By [trimeche], there exists a signed measure $\mu^{\mathcal{R}(\theta)}_{a_{1},a_{2}}$ supported over $B(0,\left\lVert a_{1}\right\rVert_{2}+\left\lVert a_{2}\right\rVert_{2})$ such that the equation in Conjecture˜1.6 is satisfied. It remains to determine whether this measure is nonnegative; more precisely, we must determine whether the symmetric version of this measure is nonnegative.
The paper [rosler_product] shows that Conjecture˜1.6 holds in the context of radially symmetric Bessel functions. Also, note that it has already been determined that the analogous conjecture for the nonsymmetric eigenfunctions of the Dunkl operators rather than the symmetric eigenfunctions is false, see [convolution_maximal].
As corollaries of the main results of this paper, we prove the following results about the weak convergences of the measures mentioned in Conjecture˜1.6. For the proofs of the corollaries, see Subsection˜10.1.
@@
Assume that $\theta\geq 0$ for all $N\geq 1$ and $\lim_{N\rightarrow\infty}\theta N=\infty$. Let $\mu_{a}$ and $\mu_{b}$ be compactly supported Borel probability measures over $\mathbb{R}$. Suppose $a(N),b(N)\in\mathbb{R}^{N}$ for $N\geq 1$ such that $\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{a(N)_{i}}{\theta N}\right)\rightarrow\mu_{a}$ and $\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{b(N)_{i}}{\theta N}\right)\rightarrow\mu_{b}$ in terms of moments as $N\rightarrow\infty$. Let $\mu$ be the free convolution of $\mu_{a}$ and $\mu_{b}$ as defined in [nica_speicher_2006]*Definition 12.1.
@@
    Assume that for all $N\geq 1$, the measure $\mu_{a(N),\,b(N)}^{A^{N-1}(\theta)}$ satisfies the conditions of Conjecture˜1.6. Then, $\operatorname{\mathbb{E}}_{a\sim\mu_{a(N),\,b(N)}^{A^{N-1}(\theta)}}\left[\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{a_{i}}{\theta N}\right)\right]$ converges to $\mu$ as $N\rightarrow\infty$ in distribution and in terms of moments.
@@
    Assume that for all $N\geq 1$, the measure $\mu_{a(N),\,b(N)}^{D^{N}(\theta)}$ satisfies the conditions of Conjecture˜1.6. Then, $\operatorname{\mathbb{E}}_{a\sim\mu_{a(N),\,b(N)}^{D^{N}(\theta)}}\left[\sum_{i=1}^{N}\frac{1}{2N}\delta\left(\frac{a_{i}}{\theta N}\right)+\frac{1}{2N}\delta\left(-\frac{a_{i}}{\theta N}\right)\right]$ converges to $\tilde{\mu}$ as $N\rightarrow\infty$ in distribution and in terms of moments.
Conjecture˜1.6 is known for $\theta\in\{\frac{1}{2},1,2\}$. By [crystalrandommat]*Proposition 2.3, the measure $\mu_{a(N),\,b(N)}^{A^{N-1}(\theta)}$ is given by the spectrum of the addition of two random matrices, one with spectrum given by $a(N)$ and the other with spectrum given by $b(N)$. When $\theta\in\{\frac{1}{2},1,2\}$, the matrix is uniformly random and self-adjoint with real, complex, and quaternion entries, respectively. Then, by [voiculescu], see also [matrix]*Theorem 1.1, we can deduce that the measure weakly converges to the free convolution.
For the type BC root system, we can deduce convergence to the rectangular free convolution introduced in [rectangular_free_convolution].
@@
Assume that $\theta_{0},\theta_{1}\geq 0$ for all $N\geq 1$, $\lim_{N\rightarrow\infty}\theta_{0}N=\infty$, and $\lim_{N\rightarrow\infty}\frac{\theta_{1}}{\theta_{0}N}\\ =c\in\mathbb{C}$. Let $\mu_{a}$ and $\mu_{b}$ be compactly supported Borel probability measures over $\mathbb{R}$. Suppose $a(N),b(N)\in\mathbb{R}^{N}$ for $N\geq 1$ such that $\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{a(N)_{i}}{\theta_{0}N}\right)\rightarrow\mu_{a}$ and $\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{b(N)_{i}}{\theta_{0}N}\right)\rightarrow\mu_{b}$ in terms of moments as $N\rightarrow\infty$.
@@
Assume that for all $N\geq 1$, the measure $\mu_{a(N),\,b(N)}^{BC^{N}(\theta_{0},\theta_{1})}$ satisfies the conditions of Conjecture˜1.6. Then, $\operatorname{\mathbb{E}}_{a\sim\mu^{BC^{N}(\theta_{0},\theta_{1})}_{a(N),\,b(N)}}\left[\sum_{i=1}^{N}\frac{1}{2N}\delta\left(\frac{a_{i}}{\theta_{0}N}\right)+\frac{1}{2N}\delta\left(-\frac{a_{i}}{\theta_{0}N}\right)\right]$ converges to $\mu$ as $N\rightarrow\infty$ in distribution and in terms of moments.
@@
Furthermore, we discuss generalizations of Corollaries˜1.7 and 1.8 in Subsection˜10.5 when $a(N)$ and $b(N)$ are randomly distributed based on exponentially decaying measures rather than deterministic. In particular, we use this framework to analyze the Dyson Brownian motion with random initial values in Example˜10.26. We establish that if the average of the scaled empirical distribution of the initial values converge in terms of moments to a Borel probability measure $\eta$ over $\mathbb{R}$, then the average of the scaled empirical distribution after $\alpha N$ time steps converges in terms of moments to a measure determined by $\eta$ and $\alpha$.
@@ 1.5. Uniform convergence of the Bessel functions
When $|\theta N|\rightarrow\infty$ and $\lim_{N\rightarrow\infty}\frac{\sum_{i=1}^{N}a(N)_{i}^{k}}{N(\theta N)^{k}}$ exists for all $k\in\mathbb{N}$, we can no longer compute the asymptotic coefficients; for example, in part (A) of Theorem˜1.1, $c_{\lambda}(N)$ has order $(\theta N)^{\ell(\lambda)}$ for $\lambda\in\Gamma$. In this case, we can still approximate the coefficients, although they will not converge.
@@
To prove that the Bessel functions are bounded, we assume that $\theta,\theta_{0},\theta_{1}\in\mathbb{R}_{\geq 0}$ so that we can apply Theorem˜1.4. An interesting direction for future research is to generalize these arguments to whenever $\theta,\theta_{0},\theta_{1}\in\mathbb{C}$ such that the corresponding Bessel functions exist.
1.6. Coefficients of terms with all odd degrees in the type D Bessel function
The type D Bessel function is a symmetric linear combination of terms with all even and all odd degrees. To analyze the coefficients of the terms with all odd degrees of these functions, we compute the asymptotics of
(1) $\prod_{j=1}^{N}\mathcal{D}_{j}(D^{N}(\theta))\prod_{i=1}^{\ell(\nu)}\sum_{j=1}^{N}\mathcal{D}_{j}(D^{N}(\theta))^{\nu_{i}}\prod_{j=1}^{N}x_{j}\prod_{i=1}^{\ell(\lambda)}\sum_{j=1}^{N}x_{j}^{\lambda_{i}}$
for $\nu,\lambda\in\Gamma_{\text{even}}$. These quantities are mentioned in part (h) of Theorem˜1.1 and are interesting because they are no longer polynomials in $N$ and $\theta$; however, they can be expressed as the product of a polynomial and an expression involving gamma functions, see Theorem˜8.16. In fact, we can consider analogous results for $A^{N-1}$ and $BC^{N}$, although terms with all odd degrees are not particularly important for analyzing the Bessel functions for these root systems so the results are not used while computing their coefficients, see Theorems˜12.1 and 12.3.
Perhaps more interesting are the asymptotics of (1) in the regime $\theta N\rightarrow c\in\mathbb{C}$ for $A^{N-1}$ and $D^{N}$ and the regime $\theta_{0}N\rightarrow c_{0}\in\mathbb{C}$, $\theta_{1}\rightarrow c_{1}\in\mathbb{C}$ for $BC^{N}$. For the corresponding leading order terms for the $A^{N-1}$, $BC^{N}$, and $D^{N}$ root systems, see Theorems 12.4, 12.5, and 8.23, respectively. Of course, we are most interested in the $D^{N}$ case since that is when summands which are multiples of $\prod_{j=1}^{N}x_{j}$ are present in the Bessel function. However, it is interesting that we can apply the methods that we develop for the $D^{N}$ case to the $A^{N-1}$ and $BC^{N}$ cases.
@@ 1.7. Applying a graded ring of operators to a graded vector field
In Sections 3, 4, and 5, we discuss the applications of a graded ring of operators to a graded vector field. The paper [dunkl_singular_poly] discusses the applications of operators to a graded vector field. We extend this notion by considering a graded ring of operators. The results that we obtain are relevant to Dunkl operators and in particular the Dunkl bilinear form, see Example˜5.19; for the definition of the Dunkl bilinear form, see Subsection˜2.1. We also study the notion of invertible operators, which is the focus of [dunkl_singular_poly]. See Section˜4 for the definition of invertibility. This framework and especially the content of Section˜4 are useful for analyzing the Bessel functions.

### v5

@@ 1.7. Integral representations of the Bessel functions
Suppose $\geq 1, N\mathcal{R}\subset\mathbb{R}^{N}$ is a finite root system, and $\theta\in\theta(\mathcal{R})$ is nonnegative. Suppose $a\in\mathbb{R}^{N}$. There exists a unique Borel probability measure $\nu^{\mathcal{R}(\theta)}_{a}$ whose support is contained in the convex hull of $H(\mathcal{R})a$ such that
$E_{a}^{\mathcal{R}}(x)=\int_{\mathbb{R}^{N}}e^{\sum_{i=1}^{N}x_{i}\epsilon_{i}}d\nu^{\mathcal{R}(\theta)}_{a}(\epsilon)$
for all $x\in\mathbb{C}^{N}$. Furthermore, $\text{supp}(\nu_{a}^{\mathcal{R}(\theta)})\cap H(\mathcal{R})a$ is nonempty and the Borel probability measure
$\nu_{a}^{\text{sym; }\mathcal{R}(\theta)}\triangleq\frac{1}{|H(\mathcal{R})|}\sum_{h\in H(\mathcal{R})}\nu_{ha}^{\mathcal{R}(\theta)}$
@@
$J^{\mathcal{R}(\theta)}_{a}(x)=\int_{\mathbb{R}^{N}}e^{\sum_{i=1}^{N}x_{i}\epsilon_{i}}d\nu_{a}^{\text{sym; }\mathcal{R}(\theta)}(\epsilon)$
@@
From [dunkloperators, dunklbessel], if $\theta\in\theta(\mathcal{R})$ such that $\theta(r)$ has nonnegative real part for all $r\in\mathcal{R}$, then $\theta\in\Theta(\mathcal{R})$, or equivalently, $\mathcal{D}(\mathcal{R}(\theta))$ is invertible. See Subsection˜2.1 for the definitions of these notions. The invertibility of the Dunkl operator implies the existence of it associated eigenfunctions, see Theorems˜2.5 and 2.8. In particular, this implies that the eigenfunctions exist in the context of Theorem˜1.3.
The intertwining measure $\nu_{a}^{\mathcal{R}(\theta)}$ is equivalent to the intertwining operator between the Dunkl operator associated to $\mathcal{R}$ and $\theta$ and $\partial$, the derivative operator which is equivalent to setting $\theta=0$. For the definition of an intertwining operator, see Definition˜3.7 and for the explanation of this equivalence, see Example˜5.18. We study intertwiners in a more general setting in Theorem˜4.5.
In addition, we can analyze the product of two Bessel functions. We state a well-known conjecture about the existence of an integral representation of such products; it can easily be extended to the product of an arbitrary number of Bessel functions. The conjecture similarly assumes that the multiplicity function is nonnegative.
@@
The measure $\mu_{a_{1},a_{2}}^{\mathcal{R}(\theta)}$ is clearly not unique, by the symmetry of $J_{a}^{\mathcal{R}(\theta)}(x)$. For simplicity, we may assume that it is symmetric. By [trimeche], there exists a distribution $\mu^{\mathcal{R}(\theta)}_{a_{1},a_{2}}$ supported over $B(0,\left\lVert a_{1}\right\rVert_{2}+\left\lVert a_{2}\right\rVert_{2})$ such that the equation in Conjecture˜1.5 is satisfied. It remains to determine whether the symmetric version of this distribution is nonnegative.
Conjecture˜1.5 is established for $A^{N-1}(\theta)$ when $\theta\in\{\frac{1}{2},1,2\}$ as well as for additional choices of the root system and multiplicity function, see the paper [rosler_product]. The paper also shows that the conjecture holds in the context of radially symmetric Bessel functions.
In addition, analogues of Conjecture˜1.5 have been established for the Schur functions, Hall-Littlewood polynomials, zonal polynomials, quaternionic zonal polynomials, and Schur’s $Q$-functions, see [macdonald]. However, the analogous conjecture for the nonsymmetric eigenfunctions of the Dunkl operators is false, see [convolution_maximal].
As corollaries of the main results of this paper, we prove the following results about the weak convergences of the measures mentioned in Conjecture˜1.5. For the proofs of the corollaries, see Subsection˜10.1. First, we define the notion of convergence in terms of power sums, which implies convergence in terms of moments.
Definition 1.6.
Suppose $\eta$ is a distribution over $\mathbb{R}$ with finite moments. For $N\geq 1$, suppose $\eta_{N}$ is a Borel probability measure over $\mathbb{R}^{N}$. Then, $\eta_{N}$ converges to $\eta$ in terms of power sums if for all $\nu\in\Gamma$,
(1) $\lim_{N\rightarrow\infty}\frac{1}{N^{\ell(\nu)}}\mathbb{E}_{a\sim\eta_{N}}\left[p_{\nu}(a)\right]=\prod_{i=1}^{\ell(\eta)}\mathbb{E}_{a\sim\eta}[a^{\nu_{i}}].$
Converging in terms of power sums is similar to requiring that $\eta_{N}\approx\otimes^{N}\eta$. Conversely, we have that if $\eta_{N}=\otimes^{N}\eta$ and $\eta$ has finite moments, then $\eta_{N}$ converges to $\eta$ in terms of power sums as $N\rightarrow\infty$. This is because for all $\nu\in\Gamma$,
$\mathbb{E}_{a\sim\eta_{N}}\left[p_{\nu}(a)\right]=\frac{N!\prod_{i=1}^{\ell(\nu)}\mathbb{E}_{a\sim\mu}[a^{\nu_{i}}]}{(N-\ell(\nu))!}+O(N^{\ell(\nu)-1}),$
since $\eta$ has finite moments. Moreover, observe that when $\eta_{N}$ is distributed as a point mass for all $N\geq 1$, convergence in terms of power sums is equivalent to convergence in terms of moments.
Converging in terms of moments corresponds to requiring that (1) is satisfied when $\nu\in\Gamma$ has length one. Another weaker condition that is often studied is to require that (1) is satisfied when $\nu\in\Gamma$ has length one or consists of two equal parts. This would imply that $\frac{1}{N^{\ell(\nu)}}p_{\nu}(\eta_{N})$ converges to $\prod_{i=1}^{\ell(\nu)}\mathbb{E}_{a\sim\eta}[a^{\nu_{i}}]$ in probability for all $\nu\in\Gamma$.
Furthermore, if $\eta$ is determined by its moments, then after applying [billingsley]*Theorem 30.2, convergence in terms of power sums implies that $\mathbb{E}_{a\sim\eta_{N}}\left[\frac{1}{N}\sum_{i=1}^{N}\delta_{(a_{i})}\right]$ converges to $\eta$ in distribution as $N\rightarrow\infty$. In particular, Carleman’s condition implies that $\eta$ is determined by its moments when it is compactly supported.
@@
Assume that $\theta\geq 0$ for all $N\geq 2$ and $\lim_{N\rightarrow\infty}\theta N=\infty$. Let $\mu_{a}$ and $\mu_{b}$ be Borel probability measures over $\mathbb{R}$ with finite moments. Suppose $a(N),b(N)\in\mathbb{R}^{N}$ for $N\geq 2$ such that $\delta_{a(N)}(x\mapsto\frac{x}{\theta N})\rightarrow\mu_{a}$ and $\delta_{b(N)}(x\mapsto\frac{x}{\theta N})\rightarrow\mu_{b}$ in terms of moments as $N\rightarrow\infty$. Let $\mu$ be the free convolution of $\mu_{a}$ and $\mu_{b}$.
@@
    Assume that for all $N\geq 2$, the measure $\mu_{a(N),\,b(N)}^{A^{N-1}(\theta)}$ satisfies the conditions of Conjecture˜1.5. Then, $\mu_{a(N),\,b(N)}^{A^{N-1}(\theta)}(x\mapsto\frac{x}{\theta N})$ converges to $\mu$ as $N\rightarrow\infty$ in terms of power sums.
@@
    Assume that for all $N\geq 2$, the measure $\mu_{a(N),\,b(N)}^{D^{N}(\theta)}$ satisfies the conditions of Conjecture˜1.5. Then, $\frac{1}{2}\mu_{a(N),\,b(N)}^{D^{N}(\theta)}(x\mapsto\frac{x}{\theta N})+\frac{1}{2}\mu_{a(N),\,b(N)}^{D^{N}(\theta)}(x\mapsto-\frac{x}{\theta N})$ converges to $\tilde{\mu}$ as $N\rightarrow\infty$ in terms of power sums.
The free convolution is introduced in [voiculescu1] and is defined for pairs of measures with finite variance in [maassen]. However, the results of this paper require $\mu_{a}$ and $\mu_{b}$ to have finite moments. For a discussion of the concepts of free probability, see [nica_speicher_2006].
If we assume that there is at most one measure with free cumulants given by adding the free cumulants of $\mu_{a}$ and $\mu_{b}$, we have that this corollary proves the existence of the free convolution as a probability measure when $\mu_{a}$ and $\mu_{b}$ have finite moments. This follows from considering when $\theta\in\{\frac{1}{2},1,2\}$ so that Conjecture˜1.5 is known to be true.
When $\theta\in\{\frac{1}{2},1,2\}$, by [crystalrandommat]*Proposition 2.3, the measure $\mu_{a(N),\,b(N)}^{A^{N-1}(\theta)}$ is given by the spectrum of the addition of two random matrices, one with spectrum given by $a(N)$ and the other with spectrum given by $b(N)$. Moreover, when $\theta=\frac{1}{2}$, $1$, and $2$, the matrix is uniformly random and self-adjoint with real, complex, and quaternion entries, respectively. Then, by [voiculescu2], see also [matrix]*Theorem 1.1, we can deduce that the measure weakly converges to the free convolution.
For the type BC root system, we deduce convergence to the rectangular free convolution introduced in [rectangular_free_convolution].
@@
Assume that $\theta_{0},\theta_{1}\geq 0$ for all $N\geq 2$, $\lim_{N\rightarrow\infty}\theta_{0}N=\infty$, and $\lim_{N\rightarrow\infty}\frac{\theta_{1}}{\theta_{0}N}\\ =c\in\mathbb{C}$. Let $\mu_{a}$ and $\mu_{b}$ be Borel probability measures over $\mathbb{R}$ with finite moments. Suppose $a(N),b(N)\in\mathbb{R}^{N}$ for $N\geq 2$ such that $\delta_{a(N)}(x\mapsto\frac{x}{\theta_{0}N})\rightarrow\mu_{a}$ and $\delta_{b(N)}(x\mapsto\frac{x}{\theta_{0}N})\rightarrow\mu_{b}$ in terms of moments as $N\rightarrow\infty$.
@@
Assume that for all $N\geq 2$, the measure $\mu_{a(N),\,b(N)}^{BC^{N}(\theta_{0},\theta_{1})}$ satisfies the conditions of Conjecture˜1.5. Then, $\frac{1}{2}\mu_{a(N),\,b(N)}^{BC^{N}(\theta_{0},\theta_{1})}(x\mapsto\frac{x}{\theta_{0}N})+\frac{1}{2}\mu_{a(N),\,b(N)}^{BC^{N}(\theta_{0},\theta_{1})}(x\mapsto-\frac{x}{\theta_{0}N})$ converges to $\mu$ as $N\rightarrow\infty$ in terms of power sums.
@@
Furthermore, we discuss generalizations of Corollaries˜1.7 and 1.8 in Subsection˜10.5 when $a(N)$ and $b(N)$ are randomly distributed based on exponentially decaying measures rather than deterministic. In particular, we use this framework to analyze the Dyson Brownian motion (DBM) with random initial values in Example˜10.28. We establish that if the average of the scaled empirical distribution of the initial values converge in terms of power sums to a Borel probability measure $\eta$ over $\mathbb{R}$, then the average of the scaled empirical distribution after $\alpha N$ time steps converges in terms of power sums to a measure determined by $\eta$ and $\alpha$.
@@ 1.8. Uniform convergence of the Bessel functions
When $|\theta N|\rightarrow\infty$ and $\lim_{N\rightarrow\infty}\frac{\sum_{i=1}^{N}a(N)_{i}^{k}}{N(\theta N)^{k}}$ exists for all $k\in\mathbb{N}$, we can no longer compute the asymptotic coefficients; for example, in part (A) of Theorem˜1.2, $c_{\lambda}(N)$ has order $(\theta N)^{\ell(\lambda)}$ for $\lambda\in\Gamma$. In this case, we can still approximate the coefficients, although they will not converge.
@@
To prove that the Bessel functions are bounded, we assume that $\theta,\theta_{0},\theta_{1}\in\mathbb{R}_{\geq 0}$ so that we can apply Theorem˜1.3. An interesting direction for future research is to generalize these arguments to whenever $\theta,\theta_{0},\theta_{1}\in\mathbb{C}$ such that the corresponding Bessel functions exist.
@@ 1.9. Applying a graded ring of operators to a graded vector field
In Sections 3, 4, and 5, we discuss the applications of a graded ring of operators to a graded vector field. The paper [dunkl_singular_poly] discusses the applications of operators to a graded vector field. We extend this notion by considering a graded ring of operators. The results that we obtain are relevant to Dunkl operators and in particular the Dunkl bilinear form, see Example˜5.18; for the definition of the Dunkl bilinear form, see Subsection˜2.1.
We also study the notion of invertible operators, which is the focus of [dunkl_singular_poly]; see Definition˜4.1 for the definition of invertibility. In particular, we prove Theorem˜4.5, which establishes many equivalent conditions for the invertibility of an operator. By the equivalence of conditions (a) and (i) of the theorem, for a fixed multiplicity function, the Bessel function exists if and only if the Dunkl operator is invertible. Due to this, the content of Section˜4 is useful for analyzing the Bessel functions. In particular, in Section˜9, we obtain the asymptotic coefficients of the Bessel functions by computing the inverses of the matrices that store the values of the Dunkl bilinear form.